# High-dimensional analysis of ridge regression for non-identically distributed data with a variance profile

### v3

@@ 1 Introduction
High-dimensionality is a subject of interest in the field of statistics, especially in regression problems, driven by the advent of massive data sets. This context gives rise to unexpected phenomena and contradictions with established statistical heuristics when the dimension $p$ of the predictors is fixed and the number $n$ of observations tends to infinity. These phenomena particularly appear in the context of linear regression. Indeed, as the sample size and dimension of acquired data increase, the study of this model is different from the classical framework. In the asymptotic regime where $\min(n,p)\to+\infty$ and $\frac{p}{n}\to c>0$, one can notably mention the occurrence of the double descent phenomenon corresponding to estimators that both interpolates the data and show good generalization performances [BHMM19]. In this asymptotic setting, using tools from random matrix theory (RMT), many authors have therefore focused on the consequences of high-dimensionality on linear regression, see e.g. [DW18, Bac24, HMRT22, LC18] and references therein. In this paper, we focus on the linear regression model
@@
where $x\in\mathbb{R}^{p}$ is a vector of random predictors, $\varepsilon\in{\mathbb{R}}$ is a noise vector independent of $x$ with $\mathbb{E}[\varepsilon]=0$ and $\mathbb{E}[\varepsilon^{2}]=\sigma^{2}>0$, $\beta_{\ast}\in{\mathbb{R}}^{p}$ is a vector of unknown parameters, and $y\in{\mathbb{R}}$ is the observed response. Let consider a training sample $(x_{1},y_{1}),...,(x_{n},y_{n})$ following this linear model, that is for $i=1,...,n$, one has that $y_{i}=x_{i}^{\top}\beta_{\ast}+\varepsilon_{i}$ where the noise term $\varepsilon_{i}$ is defined as above. Then one obtain from (1.1) that $Y_{n}=X_{n}\beta_{\ast}+\varepsilon_{n},$ where $X_{n}=(x_{1}|\dots|x_{n})^{\top}\in\mathbb{R}^{n\times p}$, $Y_{n}=(y_{1}|\dots|y_{n})^{\top}\in\mathbb{R}^{n}$ and $\varepsilon_{n}=(\varepsilon_{1}|\dots|\varepsilon_{n})^{\top}\in\mathbb{R}^{n}$.
Classically, the predictors are assumed to be independent and identically distributed (iid) data, meaning that the rows of the matrix $X_{n}$ are independent vectors sampled from the same probability distribution. In this paper, we propose to depart from this assumption by considering the setting where the rows of $X_{n}$ are independent but non-identically distributed. To this end, we suppose that $X_{n}$ is expressed in the following form $X_{n}=\Upsilon_{n}\circ Z_{n},$ where $\circ$ denotes the Hadamard product between two matrices, $Z_{n}=(Z_{ij})$ has iid centered entries with variance one, and $\Upsilon_{n}=(\gamma_{ij}^{(n)})$ is a deterministic matrix. To simplify the notation, we shall sometimes write $\gamma_{ij}^{(n)}=\gamma_{ij}$ and thus drop the (possibly) dependence of $\gamma_{ij}$ on $n$. The matrix $\Gamma_{n}=(\gamma_{ij}^{2})\in{\mathbb{R}}^{n\times p}$ governs the variance of the entries of $X_{n}$, and it is called a variance profile.
@@
The main goal of this paper is then to understand how assuming such a variance profile for $X_{n}$ influences the statistical properties of ridge regression in the linear model (1.1) when compared to the standard assumption of iid observations. In this setting, our approach also allows to analyze the performances of the minimum norm least-squares estimator when the ridge regularization parameter goes to zero.
As an application, our methodology offers a novel tool for analyzing data that arises from a mixture model that is a classical framework in machine learning, when the data are sampled from multiple underlying subpopulations or classes. Indeed, consider latent class variables $C_{1},\ldots,C_{K}$, which determine the class membership of each feature vector $x_{i}$. Formally, for each $i$ and each class $k\in\{1,\ldots,K\}$, we assume that the latent class variable $C_{i}$ follows a categorical distribution: $\forall 1\leq i\leq n,\quad\mathbb{P}(C_{i}=k)=\pi_{k},$ Within each class, the random predictor $x_{i}$ is then assumed to follow a specific covariance structure. Conditional on $\{C_{i}=k\}$, we model the predictors as $x_{i}=S_{k}^{1/2}x_{i}^{\prime},$ where $S_{k}=\text{diag}(s_{k,1}^{2},\ldots,s_{k,p}^{2})$ is a diagonal matrix that characterizes the covariance structure of the predictors within class $k$. The variances $s_{k,j}^{2}$ of the predictors vary across classes, reflecting potential heterogeneity in the data. Then, given the class labels $C_{1},\ldots,C_{n}$, the resulting matrix of predictors $X_{n}$ exhibits a variance profile governed by the matrix $\Gamma_{n}=(s_{C_{i},j}^{2})\in\mathbb{R}^{n\times p},$ where each entry $s_{C_{i},j}^{2}$ corresponds to the variance of the $j$-th feature for the $i$-th feature vector, determined by its class membership $C_{i}$. Hence, our approach allows to investigate the double descent phenomenon in the context of high-dimensional ridge regression when the data follow a mixture model, an aspect that has not been previously explored in the literature, to the best of our knowledge.
@@
Our analysis also includes the study of the minimum least-square estimator defined as
$\hat{\theta}=\operatorname*{arg\,min}_{\theta\in{\mathbb{R}}^{p}}\left\{\|\theta\|\;:\;\theta\mbox{ minimizes }\frac{1}{n}\|Y_{n}-X_{n}\theta\|^{2}\right\},$
to which the ridge regression estimator converges when $\lambda$ tends to zero. The estimator $\hat{\theta}$ is also known to be the solution found by gradient descent when initialized to zero, see e.g. [HMRT22][Proposition 1].
To study the statistical performances of the ridge regression estimator, we analyse its predictive risk, denoted $\hat{r}_{\lambda}^{test}(X_{n})$, and its train risk, denoted $\hat{r}_{\lambda}^{train}(X_{n})$, defined as
$\hat{r}_{\lambda}^{test}(X_{n})=\mathbb{E}[(\tilde{y}-\tilde{x}^{\top}\hat{\theta}_{\lambda})^{2}|X_{n}],$ (1.3)
$\hat{r}_{\lambda}^{train}(X_{n})=\frac{1}{n}\mathbb{E}[\parallel Y_{n}-X_{n}\hat{\theta}_{\lambda}\parallel^{2}|X_{n}],$ (1.4)
where $(\tilde{y},\tilde{x})\in{\mathbb{R}}\times{\mathbb{R}}^{p}$ is independent from $(Y_{n},X_{n})$ and satisfies
$\tilde{y}=\tilde{x}^{\top}\beta_{\ast}+\tilde{\varepsilon},\mbox{ with }\mathbb{E}[\tilde{\varepsilon}]=0,\;\mathbb{E}[\tilde{\varepsilon}^{2}]=\sigma^{2}\ \mathrm{and}\ \widetilde{X},\ \tilde{\varepsilon}\ \mathrm{independent}.$
In the above formula, $\tilde{x}=\widetilde{S}_{p}^{1/2}\widetilde{z}$ with $\widetilde{z}\in{\mathbb{R}}^{p}$ a random vector with iid centered entries and variance one, and $\widetilde{S}_{p}=\mathbb{E}[\tilde{x}\tilde{x}^{\top}]=\mathop{\rm diag}\nolimits(\tilde{\gamma}_{1}^{2},\ldots,\tilde{\gamma}_{p}^{2})$ denotes the variance profile of $\tilde{x}$. Note that the risk $\hat{r}_{\lambda}^{test}(X_{n})$ is conditioned on the predictors $X_{n}$, and it is thus a random variable.
@@
Assumption 1.4.
Assumption 1.5.
The coefficient $\alpha>0$ represents the average amount of signal strength in model (1.1). Under Assumption 1.1, the expectation in (1.3) used to define the predictive risk is thus taken with respect to both the randomness of the vector of coefficients $\beta_{\ast}$, the vector $\tilde{x}$ and the additive noise $\varepsilon_{n}$ and $\tilde{\varepsilon}$. Assumptions 1.2 to 1.5 ensures that the spectrum of $\frac{1}{n}X_{n}^{\top}X_{n}$ well behave in the high dimensional framework, that is whenever $\min(n,p)\to+\infty$ and $\frac{p}{n}\to c>0$. Indeed, it is proved in [HLN07] that Assumptions 1.2 and 1.3 guarantees that the empirical singular value distribution of $X_{n}$ converges, in the high dimensional regime, to a distribution solution of a fixed point equation. Note that these assumptions are not limited to the case of random Gaussian data. On the other hand, [AEK17b] proves that Assumptions 1.4 and 1.5 ensures that the spectrum of $\frac{1}{n}X_{n}^{\top}X_{n}$ is bounded away from 00. Note that we only use this condition to study the behavior of the risks in the ridge (less) framework, that is whenever $\lambda=0$. Hence Assumptions 1.4 and 1.5 can be ignored as long as $\lambda\neq 0$.
@@ 1.1 Main contributions
Recall that the estimation of $Y_{n}$ by ridge regression is
$\hat{Y}_{\lambda}=X_{n}\hat{\theta}_{\lambda}=A_{\lambda}Y_{n},\quad\mbox{where}\quad A_{\lambda}=X_{n}(X^{\top}_{n}X_{n}+n\lambda I_{p})^{-1}X_{n}^{\top}.$
Then, the degrees of freedom (DOF) of the estimator $\hat{\theta}_{\lambda}$, that is defined as
$\widehat{df}_{1}(\lambda)=\operatorname{Tr}[A_{\lambda}]=\operatorname{Tr}[\widehat{\Sigma}_{n}(\widehat{\Sigma}_{n}+\lambda I_{p})^{-1}],\mbox{ where }\widehat{\Sigma}_{n}=\frac{1}{n}X_{n}^{\top}X_{n},$
@@
$\widehat{df}_{1}(\lambda)\sim df_{1}(\lambda),\quad\mbox{where}\quad df_{1}(\lambda)=\operatorname{Tr}[\Sigma_{n}(\Sigma_{n}+\kappa(\lambda))^{-1}],$ (1.5)
where $\Sigma_{n}=\mathbb{E}[\widehat{\Sigma}_{n}]=\frac{1}{n}\mathop{\rm diag}\nolimits\left(\sum_{i=1}^{n}\gamma^{2}_{i1},\ldots,\sum_{i=1}^{n}\gamma^{2}_{ip}\right),$ and $\kappa(\lambda)$ is diagonal matrix that depends upon the regularization parameter $\lambda$ and the variance profile matrix $\Gamma_{n}$.
Throughout the paper, the meaning of the equivalence notation $A_{n}\sim B_{n}$ between two random variables is $\lim_{n\to\infty,\;p/n\to c}|A_{n}-B_{n}|=0,\;\mbox{almost surely}.$
@@
Hence, the equivalence relation (1.5) indicates that the DOF of the ridge regression estimator for the empirical covariance matrix $\widehat{\Sigma}_{n}$ corresponds to the DOF computed with its expected version $\Sigma_{n}$ (the usual population covariance matrix for iid data), and another additive regularization structure than $\lambda I_{p}$ that is given by the diagonal matrix $\kappa(\lambda)$ whose explicit expression is given in Section 3.
Then, the second and main contribution of the paper is to derive a deterministic equivalent of the predictive risk $\hat{r}_{\lambda}^{test}(X_{n})$ and the training risk $\hat{r}_{\lambda}^{train}(X_{n})$ in the case where the number of samples $n$ and the dimension $p$ tend to infinity at a proportional rate, that is $\lim\limits_{n\to+\infty}\frac{p}{n}=c>0$. This deterministic equivalent allows us to understand the influence of the ratio $c$ on the predictive risk and to also analyze the effect of the signal strength $\alpha$. We also study the convergence of the predictive risk as $\lambda$ tends to 00 to analyze the statistical properties the minimum norm least square estimator. In this setting, it appears a phenomenon arising from the curse of dimensionality that is commonly known as double descent for iid data. This phenomenon contradicts the consensus heuristic that, when a model becomes over-parameterized, then the predictive risk increases due to overfitting of the training data and the model is no longer capable of generalizing. This double descent has been thoroughly studied in the case of high-dimensional linear regression using tool from RMT, see e.g. [HMRT22, Bac24, BHX20] and references therein. In this paper, we show that it also occurs for non iid data with a variance profile. Our deterministic equivalent of the predictive risk also allows to derive the asymptotic behavior of an optimal choice for the regularization parameter and to compare it to the one obtained for iid data in [DW18].
As a third contribution, using synthetic data and various illustrative examples of variance profile, we conduct numerical experiments to verify the accuracy of our deterministic equivalent of the predictive risk using finite samples.
@@
We also investigate the similarities and differences that exist between the standard setting of iid data and the one of non-identically distributed data with a variance profile. For example, if the variance profile is assumed to be quasi doubly stochastic in the sense that
@@
then, we prove in this paper that
$\lim_{n\to\infty,\;p/n\to c}\hat{r}_{0}^{test}(X_{n})=\left\{\begin{array}[]{ccc}\sigma^{2}\frac{c}{1-c}+\sigma^{2}&\mbox{if}&c<1\\ \alpha^{2}\left(1-\frac{1}{c}\right)+\sigma^{2}\frac{1}{c-1}&\mbox{if}&c>1\end{array}\right.,\;\mbox{almost surely}.$ (1.7)
The above result corresponds to the known asymptotic limit of the predictive risk of the minimum norm least squares estimator for iid data when the entries of $X_{n}$ are independent centered random variables with variance $1$, see e.g. [HMRT22], which is referred to as a constant variance profile in this paper (that is $\gamma_{ij}=1$). In this setting, the predictive risk of $\hat{\theta}$ is thus increasing with $p$ up to $p/n<1$ and then decreasing for $p>n$ (that is beyond the interpolation threshold $p=n$) which is classically referred to as the double descent phenomenon as illustrated in Figure 1(a).
By going beyond the quasi doubly stochastic assumption (1.7), the results of this paper, on the asymptotic limit of the predictive risk, also allow to exhibit variance profiles for which the predictive risk has a shape that differs from double descent. This is illustrated in Figure 1(b), where the deterministic equivalent of the predictive risk has a triple descent behavior (as a function of $p/n$) for the following piecewise constant variance profile

### v4

@@ 1 Introduction
High-dimensionality is a subject of interest in the field of statistics, especially in regression problems, driven by the advent of massive data sets. This context gives rise to unexpected phenomena and contradictions with established statistical heuristics when the dimension $p$ of the predictors is fixed and the number $n$ of observations tends to infinity. These phenomena particularly appear in the context of linear regression. Indeed, as the sample size and dimension of acquired data increase, the study of this model is different from the classical framework. In the asymptotic regime where $\min(n,p)\to+\infty$ and $\frac{p}{n}\to c>0$, one can notably mention the occurrence of the double descent phenomenon corresponding to estimators that both interpolates the data and show good generalization performances [BHMM19]. This phenomenon contradicts the consensus heuristic that, when a model becomes over-parameterized, then the predictive risk increases due to overfitting of the training data and the model is no longer capable of generalizing. This double descent phenomenon has been thoroughly studied in the case of high-dimensional linear regression with independent and identically distributed (iid) data using tool from RMT, see e.g. [HMRT22, Bac24, BHX20] and references therein. For instance, when the entries of $X_{n}$ are independent centered random variables with variance 1, the empirical prediction risk $\hat{r}_{0}^{\text{test}}(X_{n})$ (which we will define later) admits the following expression [HMRT22] when $n\to\infty$ and $p/n\to c$:
$\lim_{n\to\infty,\;p/n\to c}\hat{r}_{0}^{\text{test}}(X_{n})=\begin{cases}\sigma^{2}\dfrac{c}{1-c}+\sigma^{2}\mbox{ if }c<1,\\ \alpha^{2}\left(1-\dfrac{1}{c}\right)+\sigma^{2}\dfrac{1}{c-1}\mbox{ if }c>1,\end{cases}\quad\text{almost surely}.$ (1.1)
This result shows that the prediction risk increases with $p$ as long as $p/n<1$, reaches a peak around the interpolation threshold $p=n$, and then decreases when $p>n$. This non-monotonic behavior is classically referred to as the double descent phenomenon and is illustrated in Figure 1(a).
In this asymptotic setting, using tools from random matrix theory (RMT), many authors have therefore focused on the consequences of high-dimensionality on linear regression, see e.g. [DW18, Bac24, HMRT22, LC18] and references therein. In this paper, we focus on the linear regression model
@@
where $x\in\mathbb{R}^{p}$ is a vector of random predictors, $\varepsilon\in{\mathbb{R}}$ is a noise vector independent of $x$ with $\mathbb{E}[\varepsilon]=0$ and $\mathbb{E}[\varepsilon^{2}]=\sigma^{2}>0$, $\beta_{\ast}\in{\mathbb{R}}^{p}$ is a vector of unknown parameters, and $y\in{\mathbb{R}}$ is the observed response. Let consider a training sample $(x_{1},y_{1}),...,(x_{n},y_{n})$ following this linear model, that is for i=1,,ni=1,...,n, one has that $y_{i}=x_{i}^{\top}\beta_{\ast}+\varepsilon_{i}$ where the noise term $\varepsilon_{i}$ is defined as above. Then one obtain from Equation (1.2) that $Y_{n}=X_{n}\beta_{\ast}+\varepsilon_{n},$ where $X_{n}=(x_{1}|\dots|x_{n})^{\top}\in\mathbb{R}^{n\times p}$, $Y_{n}=(y_{1}|\dots|y_{n})^{\top}\in\mathbb{R}^{n}$ and $\varepsilon_{n}=(\varepsilon_{1}|\dots|\varepsilon_{n})^{\top}\in\mathbb{R}^{n}$.
Classically, the predictors are assumed to be iid data, meaning that the rows of the matrix $X_{n}$ are independent vectors sampled from the same probability distribution. In this paper, we propose to depart from this assumption by considering the setting where the rows of $X_{n}$ are independent but non-identically distributed. To this end, we suppose that $X_{n}$ is expressed in the following form $X_{n}=\Upsilon_{n}\circ Z_{n},$ where $\circ$ denotes the Hadamard product between two matrices, $Z_{n}=(Z_{ij})$ has iid centered entries with variance one, and $\Upsilon_{n}=(\gamma_{ij}^{(n)})$ is a deterministic matrix. To simplify the notation, we shall sometimes write $\gamma_{ij}^{(n)}=\gamma_{ij}$ and thus drop the (possibly) dependence of $\gamma_{ij}$ on $n$. The matrix $\Gamma_{n}=(\gamma_{ij}^{2})\in{\mathbb{R}}^{n\times p}$ governs the variance of the entries of $X_{n}$, and it is called a variance profile.
@@
The main goal of this paper is then to understand how assuming such a variance profile for $X_{n}$ influences the statistical properties of ridge regression in the linear model Equation (1.2) when compared to the standard assumption of iid observations. In this setting, our approach also allows to analyze the performances of the minimum norm least-squares estimator when the ridge regularization parameter goes to zero.
The deterministic equivalents derived in this paper are not only of interest in the RMT context, as they also provide approximations of key quantities to statistics in high dimensions, namely the prediction error and the effective model complexity measured by its degrees of freedom. In the classical iid setting, such equivalents are typically used as a means to an end, enabling one to highlight regimes where ridge regression generalizes well or poorly, characterize optimal regularization parameter, and study ridge regression across the under and over-parameterized regimes. A main objective of the present work is to provide analogous consequences in the non-iid setting induced by a variance profile, and to identify which phenomena remain universal (e.g., the derivation of an optimal regularisation parameter in ridge regression) from those that depend on the variance profile (e.g., the emergence of a multiple-descent phenomenon beyond the classical double descent one).
As an application, our methodology also offers a novel tool for analyzing data that arises from a mixture model that is a classical framework in machine learning, when the data are sampled from multiple underlying subpopulations or classes. Indeed, consider latent class variables $C_{1},\ldots,C_{K}$, which determine the class membership of each feature vector $x_{i}$. Formally, for each $i$ and each class $k\in\{1,\ldots,K\}$, we assume that the latent class variable $C_{i}$ follows a categorical distribution: $\forall 1\leq i\leq n,\quad\mathbb{P}(C_{i}=k)=\pi_{k},$ Within each class, the random predictor $x_{i}$ is then assumed to follow a specific covariance structure. Conditional on $\{C_{i}=k\}$, we model the predictors as $x_{i}=S_{k}^{1/2}x_{i}^{\prime},$ where $S_{k}=\text{diag}(s_{k,1}^{2},\ldots,s_{k,p}^{2})$ is a diagonal matrix that characterizes the covariance structure of the predictors within class $k$. The variances $s_{k,j}^{2}$ of the predictors vary across classes, reflecting potential heterogeneity in the data. Then, given the class labels $C_{1},\ldots,C_{n}$, the resulting matrix of predictors $X_{n}$ exhibits a variance profile governed by the matrix $\Gamma_{n}=(s_{C_{i},j}^{2})\in\mathbb{R}^{n\times p},$ where each entry $s_{C_{i},j}^{2}$ corresponds to the variance of the $j$-th feature for the $i$-th feature vector, determined by its class membership $C_{i}$. Hence, our approach allows to investigate the double descent phenomenon in the context of high-dimensional ridge regression when the data follow a mixture model, an aspect that has not been previously explored in the literature, to the best of our knowledge.
@@
Our analysis also includes the study of the minimum least-square estimator defined as $\hat{\theta}=\operatorname*{arg\,min}_{\theta\in{\mathbb{R}}^{p}}\left\{\|\theta\|\;:\;\theta\mbox{ minimizes }\frac{1}{n}\|Y_{n}-X_{n}\theta\|^{2}\right\},$ to which the ridge regression estimator converges when $\lambda$ tends to zero. The estimator $\hat{\theta}$ is also known to be the solution found by gradient descent when initialized to zero, see e.g. [HMRT22][Proposition 1].
To study the statistical performances of the ridge regression estimator, we analyse its empirical predictive risk, denoted $\hat{r}_{\lambda}^{test}(X_{n})$, and its train risk, denoted $\hat{r}_{\lambda}^{train}(X_{n})$, defined as
$\hat{r}_{\lambda}^{test}(X_{n})=\mathbb{E}[(\tilde{y}-\tilde{x}^{\top}\hat{\theta}_{\lambda})^{2}|X_{n}],\quad\mbox{ and }\quad\hat{r}_{\lambda}^{train}(X_{n})=\frac{1}{n}\mathbb{E}[\parallel Y_{n}-X_{n}\hat{\theta}_{\lambda}\parallel^{2}|X_{n}],$ (1.4)
where $(\tilde{y},\tilde{x})\in{\mathbb{R}}\times{\mathbb{R}}^{p}$ is independent from $(Y_{n},X_{n})$ and satisfies $\tilde{y}=\tilde{x}^{\top}\beta_{\ast}+\tilde{\varepsilon},\mbox{ with }\mathbb{E}[\tilde{\varepsilon}]=0,\;\mathbb{E}[\tilde{\varepsilon}^{2}]=\sigma^{2}\ \mathrm{and}\ {\color[rgb]{0,0,0}\widetilde{x}},\ \tilde{\varepsilon}\ \mathrm{independent}.$ In the above formula, $\tilde{x}=\widetilde{S}_{p}^{1/2}\widetilde{z}$ with $\widetilde{z}\in{\mathbb{R}}^{p}$ a random vector with iid centered entries and variance one, and $\widetilde{S}_{p}=\mathbb{E}[\tilde{x}\tilde{x}^{\top}]=\mathop{\rm diag}\nolimits(\tilde{\gamma}_{1}^{2},\ldots,\tilde{\gamma}_{p}^{2})$ denotes the variance profile of $\tilde{x}$. Note that the risk $\hat{r}_{\lambda}^{test}(X_{n})$ is conditioned on the predictors $X_{n}$, and it is thus a random variable.
@@
The coefficient $\alpha>0$ represents the average amount of signal strength in model Equation (1.2). Suppose that Assumption 1.1 holds true, then the expectation in Equation (1.4) used to define the predictive risk is thus taken with respect to both the randomness of the vector of coefficients $\beta_{\ast}$, the vector $\tilde{x}$ and the additive noise $\varepsilon_{n}$ and $\tilde{\varepsilon}$. Assuming random regression coefficients $\beta_{\ast}$ allows for a tractable theoretical characterization of the training and predictive risks and it captures essential aspects of high-dimensional behavior that are otherwise more difficult to analyze. Assumption 1.2 with the supplementary Assumption 3.1 and Assumption 3.2 (to be given later on) ensure that the spectrum of $\frac{1}{n}X_{n}^{\top}X_{n}$ well behave in the high dimensional regime, that is whenever $\min(n,p)\to+\infty$ and $\frac{p}{n}\to c>0$. Indeed, it is proved in [HLN07] that Assumptions 1.2 and 1.3 guarantee that the empirical singular value distribution of $X_{n}$ converges, in the high dimensional regime, to a distribution solution of a fixed point equation. Note that these assumptions are not limited to the case of random Gaussian data.
@@ 1.1 Main contributions
Recall that the estimation of $Y_{n}$ by ridge regression is $\hat{Y}_{\lambda}=X_{n}\hat{\theta}_{\lambda}=A_{\lambda}Y_{n},\quad\mbox{where}\quad A_{\lambda}=X_{n}(X^{\top}_{n}X_{n}+n\lambda I_{p})^{-1}X_{n}^{\top}.$ Then, the degrees of freedom (DOF) of the estimator $\hat{\theta}_{\lambda}$, that is defined as
$\widehat{df}_{1}(\lambda)=\frac{1}{p}\operatorname{Tr}[A_{\lambda}]=\frac{1}{p}\operatorname{Tr}[\widehat{\Sigma}_{n}(\widehat{\Sigma}_{n}+\lambda I_{p})^{-1}],\mbox{ where }\widehat{\Sigma}_{n}=\frac{1}{n}X_{n}^{\top}X_{n},$
@@
$\widehat{df}_{1}(\lambda)\sim df_{1}(\lambda),\quad\mbox{where}\quad df_{1}(\lambda)=\frac{1}{p}\operatorname{Tr}[\Sigma_{n}(\Sigma_{n}+\kappa(\lambda))^{-1}],$ (1.5)
@@
where $\Sigma_{n}=\mathbb{E}[\widehat{\Sigma}_{n}]=\frac{1}{n}\mathop{\rm diag}\nolimits\left(\sum_{i=1}^{n}\gamma^{2}_{i1},\ldots,\sum_{i=1}^{n}\gamma^{2}_{ip}\right),$ and $\kappa(\lambda)$ is diagonal matrix that depends upon the regularization parameter $\lambda$ and the variance profile matrix $\Gamma_{n}$. Throughout the paper, the meaning of the equivalence notation $A_{n}\sim B_{n}$ between two random variables is $\lim_{n\to\infty,\;p/n\to c}|A_{n}-B_{n}|=0,\;\mbox{almost surely}.$ Hence, the equivalence relation Equation (1.5) indicates that the DOF of the ridge regression estimator for the empirical covariance matrix $\widehat{\Sigma}_{n}$ corresponds to the DOF computed with its expected version $\Sigma_{n}$ (the usual population covariance matrix for iid data), and another additive regularization structure than $\lambda I_{p}$ that is given by the diagonal matrix $\kappa(\lambda)$ whose explicit expression is given in Section 3.
The notion of degrees of freedom (DOF) plays a central role in the selection of the regularization parameter $\lambda$. In principle, $\lambda$ can be optimally selected by minimizing the predictive risk, as done in Section 3. However, this approach typically requires access to independent test data or assumes the use of underlying parameters such as the signal strength $\alpha$ which is typically unknown. To circumvent these limitations, a widely used data-driven alternative is the Generalized Cross-Validation (GCV) criterion [GHW79]. The GCV provides an approximation of the out-of-sample prediction error using only training data, without requiring an explicit data-splitting procedure. Following [GHW79], we define in this paper the GCV criterion as
$\color[rgb]{0,0,0}\hat{G}(\lambda)=\frac{\hat{r}_{\lambda}^{\text{train}}(X_{n})}{\left(1-\widehat{df}_{1}(\lambda)\right)^{2}}$}. (1.6)
Minimizing the GCV thus amounts to selecting the value of $\lambda$ that achieves an optimal balance between goodness-of-fit, as captured by the training risk, and model complexity, as measured by the degrees of freedom. In Section 3, we derive a deterministic equivalent of $\hat{G}(\lambda)$ and we discuss its minimizers through numerical experiments.
Then, the second and main contribution of the paper is to derive deterministic equivalents for both the predictive risk $\hat{r}_{\lambda}^{\text{test}}(X_{n})$ and the training risk $\hat{r}_{\lambda}^{\text{train}}(X_{n})$ in the high-dimensional regime. These deterministic equivalents allow to quantify the influence of the aspect ratio $c=\lim_{n\to+\infty}p/n$ on generalization performance, as well as to analyze the impact of the signal strength $\alpha$. In particular, we show the existence of an optimal value $\lambda_{\ast}=\frac{\sigma^{2}p}{\alpha^{2}n}$ that minimizes both the empirical predictive risk and its deterministic equivalent for any variance profile (see Corollary 3.1), and that equals the one obtained for iid data in [DW18]. This result reveals the robustness of optimal shrinkage with respect to heteroscedasticity in the predictors, and shows that the optimal scaling of ridge regularization is universal across a broad class of variance profiles.
The derivation of these deterministic equivalents relies on a precise control of resolvent quantities associated with the empirical covariance matrix $\widehat{\Sigma}_{n}$. More precisely, the expressions of both the training and predictive risks involve the diagonal entries of the resolvent $Q_{p}(z)=(\widehat{\Sigma}_{n}-zI_{p})^{-1}$ and of its derivative. As a consequence, the analysis requires establishing a strong deterministic equivalent for the diagonal of the resolvent $\Delta[Q_{p}(z)]$, rather than its normalized trace. We obtain such a result by extending and strengthening the resolvent analysis developed in [HLN07] for random matrices with a variance profile, thereby overcoming the lack of unitary invariance inherent to the non-identically distributed setting. In addition, we study the behavior of the predictive risk in the ridgeless limit $\lambda\to 0$ in order to analyze the statistical properties of the minimum-norm least-squares estimator. This regime is particularly delicate, as it is governed by the behavior of the spectrum of the population and sample covariance matrices near zero. A sharp local analysis of the spectrum around the origin is therefore required to control the limiting risk and to capture interpolation phenomena. Indeed, the behavior of the smallest non-zero eigenvalue of $\widehat{\Sigma}_{n}$ is closely linked to risk blow-ups in the double descent phenomenon. To this end, we rely on the precise spectral results established in [AEK17b], which provide the necessary information on the low-end behavior of the spectrum and allow us to rigorously characterize the asymptotic predictive risk in the interpolating regime.
@@
As a consequence of these results, we investigate the similarities and differences that exist between the standard setting of iid data and the one of non-identically distributed data with a variance profile. For example, if the variance profile is assumed to be quasi doubly stochastic in the sense that
@@
then, we prove in this paper that $\hat{r}_{0}^{test}(X_{n})$ satisfies Equation Equation (1.1) That is, the quasi doubly stochastic case corresponds to the known asymptotic limit of the predictive risk of the minimum norm least squares estimator for iid data [HMRT22] when the entries of $X_{n}$ are independent centered random variables with variance $1$ which is referred to as a constant variance profile in this paper (that is $\gamma_{ij}=1$). Consequently, the double descent phenomenon still arises for non-iid data as illustrated in Figure 1(a).
As a third contribution by going beyond the quasi doubly stochastic assumption Equation (1.1), the results of this paper also allow to exhibit variance profiles for which the predictive risk has a shape that differs from double descent. This is illustrated in Figure 1(b), where the deterministic equivalent of the predictive risk has a triple descent behavior (as a function of $p/n$) for the following piecewise constant variance profile